# SAR3D: Autoregressive 3D Object Generation and Understanding via Multi-scale 3D VQVAE

### v2

@@ 4.3 SAR3D-LLM for 3D Object Understanding
Since our 3D VQVAE model provides a comprehensive encoding of the given 3D object, it can be naturally extended to 3D object understanding. Following PointLLM [85], we align the latent space of our pre-trained 3D VQVAE to a large language model, e.g., LLaMA [72, 73]. As briefed in Sec. 3.2, the encoded 3D tokens $R$ are projected to the language latent space, and concatenated with the text instruction tokens $Z_{t}$. Here, we directly use the output tokens from pre-trained SAR3D VQVAE to the projector $\Gamma_{\text{proj}}$. Since we only study the 3D captioning [43] task here, the instruction tokens $Z_{t}$ is fixed to $\tilde{Z_{t}}$ which are tokenized from “Give a concise interpretation of the 3D data presented here”. The final framework, SAR3D-LLM, supports both detailed 3D captioning given 3D object and simultaneous 3D generation and captioning given text or image.
@@ 5.1 Single Image to 3D
The qualitative comparisons between SAR3D and existing methods is also included in Fig. 4. Compared to single-image to 3D methods like OpenLRM[20], and multi-view image to 3D methods like LGM [70], our approach achieves better 3D consistency across views and reduces distortion in generated 3D objects. Compared to native 3D diffusion models like LN3Diff [34], SAR3D produces more complete 3D models. Additional qualitative results are shown in Fig. 5. For more comparisons with other methods, please refer to the Supp Mat.
@@ 6 Limitations
The first limitation is that, while SAR3D can generate high-quality 3D objects and detailed interpretations, it currently relies on two separate autoregressive models. Future work could focus on developing a truly multimodal model [80] capable of processing tokens that integrate both text and 3D information, producing both 3D and text outputs. Besides, the quality of the geometry and texture is limited by volume rendering. Using more efficient 3D representions [28] or cascaded generation [93] will further boost the overall quality. Finally, although our method demonstrates inherent scalability, its scaling behavior has not been thoroughly validated here due to limited resources. We believe that with more resources, our method has the potential to demonstrate favorable scaling laws in 3D generation and understanding.
@@ 7 Conclusion
Appendix A Multi-scale quantization and interpolation
@@
Appendix B Transformer blocks
@@
Appendix C More 3D captioning results
@@
Appendix D More image-to-3D comparison
As illustrated in Fig. S3, we show more reults to compare our SAR3D with three categories of methods: single-image to 3D methods (Splatter-Image [68], OpenLRM [20, 26]), multi-view image to 3D methods (One-2-3-45 [40], Lara [9], CRM [79], LGM [70]), and native 3D diffusion models (Shap-E [29], LN3Diff-image [34]). Compared to baseline methods, our SAR3D generates intact, distortion-free results and delivers high-quality visual effects in both reference and novel views.

### v3

@@ 4.3 SAR3D-LLM for 3D Object Understanding
Since our 3D VQVAE model provides a comprehensive encoding of the given 3D object, it can be naturally extended to 3D object understanding. Following PointLLM [83], we fine-tune LLaMA [69, 70, 12] in two stages: first, we finetune the projector $\Gamma_{\text{proj}}$ while freezing the LLM $\Gamma_{\text{llm}}$ to align 3D features with the text token space; second, we fine-tune both. As briefed in Sec. 3.2, the encoded 3D tokens $R$ are projected to the language latent space, and concatenated with the text instruction tokens $Z_{t}$. Here, we directly use the output tokens from pre-trained SAR3D VQVAE to the projector $\Gamma_{\text{proj}}$. Since we only study the 3D captioning [41] task here, the instruction tokens $Z_{t}$ is fixed to $\tilde{Z_{t}}$ which are tokenized from “Give a concise interpretation of the 3D data presented here”. The final framework, SAR3D-LLM, supports both detailed 3D captioning given 3D object and simultaneous 3D generation and captioning given text or image.
@@ 5.1 Single Image to 3D
The qualitative comparisons between SAR3D and existing methods is also included in Fig. 4. Compared to single-image to 3D methods like OpenLRM[19], and multi-view image to 3D methods like LGM [67], our approach achieves better 3D consistency across views and reduces distortion in generated 3D objects. Compared to native 3D diffusion models like LN3Diff [32], SAR3D produces more complete 3D models. Additional qualitative results are shown in Fig. 5. For more comparisons with other methods and details for training data, please refer to the Supp Mat.
@@ 6 Limitations
The first limitation is that, while SAR3D can generate high-quality 3D objects and detailed interpretations, it currently relies on two separate autoregressive models. Future work could focus on developing a truly multimodal model [78] capable of processing tokens that integrate both text and 3D information, producing both 3D and text outputs. Besides, the quality of the geometry and texture is limited by volume rendering. Using more efficient 3D representations [26] or cascaded generation [90] will further boost the overall quality. Finally, our method’s scalability remains unverified due to resource limits. However, 2D scaling laws in VAR [68] suggest its feasibility. We believe that with more resources, our method has the potential to demonstrate favorable scaling laws in 3D generation and understanding.
@@ 7 Conclusion
Acknowledgements
This research is supported by the National Research Foundation, Singapore, under its NRF Fellowship Award <NRF-NRFF16-2024-0003>. This research is also supported by NTU SUG-NAP and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ Acknowledgements
Appendix A Effectiveness of 3D VQVAE
As shown in Fig. S1, our 3D Multi-scale Vector Quantized Variational Autoencoder (VQVAE) exhibits a strong capability in reconstructing complex 3D objects with diverse topologies. By leveraging a multiscale design, our model captures both global and local features, allowing it to reconstruct objects with textures, multiple holes, and diverse surface patterns.
Figure S1: Reconstruction Results of Our Multiscale VQVAE. Our VQVAE can reconstruct both complex geometries that contain multiple holes and textures.
Appendix B Multi-scale quantization and interpolation
@@
Appendix C Transformer blocks
@@
Appendix D More image-to-3D comparison
As illustrated in Fig. S3, we show more reults to compare our SAR3D with three categories of methods: single-image to 3D methods (Splatter-Image [65], OpenLRM [19, 25]), multi-view image to 3D methods (One-2-3-45 [38], Lara [8], CRM [76], LGM [67]), and native 3D diffusion models (Shap-E [27], LN3Diff-image [32]). We use pretrained models for these baseline methods, and the data statistics are provided in the Tab. S1. Compared to baseline methods, our SAR3D generates intact, distortion-free results and delivers high-quality visual effects in both reference and novel views.
Table S1: Statistics of Training Data.
Splatter-Image OpenLRM One-2-3-45 Lara CRM LGM Shap-E LN3Diff Ours
44K 580K 46K 240K 376K 80K 2M - 9M 170K 170K
Appendix E More 3D captioning results